OpenStudy (anonymous):

xy' - y = x^3 using the method of variation of parameter ???

OpenStudy (anonymous):

Do you know something about the variation of parameter method?

OpenStudy (anonymous):

it is one order diff equ i know the second order and higher so i could not do this

OpenStudy (anonymous):

@nitz and also your attachment is full of second order bro

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

i m waiting

OpenStudy (anonymous):

Nitz is not bro she is a sis..

OpenStudy (anonymous):

@waterineyes good observation :)

OpenStudy (anonymous):

Thanks bro..

OpenStudy (anonymous):

@nitz any improvements sis

OpenStudy (anonymous):

?

OpenStudy (anonymous):

can you please wait for 5 min i am solving someone other problem

OpenStudy (anonymous):

xy' - y = x^3 let y = xu then y' = u + u'x x(u + u'x) - ux = x^3 hence u + u'x - u = x^2 hence u'x = x^2 hence u' = x hence u = x^2/2 + C hence y/x = x^2/2 + C hence y = x^3/2 + Cx